### v2

# HiRAG: Retrieval-Augmented Generation with Hierarchical Knowledge

Abstract
Graph-based Retrieval-Augmented Generation (RAG) methods have significantly enhanced the performance of large language models (LLMs) in domain-specific tasks. However, existing RAG methods do not adequately utilize the naturally inherent hierarchical knowledge in human cognition, which limits the capabilities of RAG systems. In this paper, we introduce a new RAG approach, called HiRAG, which utilizes hierarchical knowledge to enhance the semantic understanding and structure capturing capabilities of RAG systems in the indexing and retrieval processes. Our extensive experiments demonstrate that HiRAG achieves significant performance improvements over the state-of-the-art baseline methods.111 https://github.com/hhy-huang/HiRAG
1 Introduction
Figure 1: The challenges faced by existing RAG systems: (1) Distant structural relationship between semantically similar entities. (2) Knowledge gap between local and global knowledge.
Retrieval Augmented Generation (RAG) Gao et al. (2023) Lewis et al. (2020) Fan et al. (2024) has been introduced to enhance the capabilities of LLMs in domain-specific or knowledge-intensive tasks. Naive RAG methods retrieve text chunks that are relevant to a query, which serve as references for LLMs to generate responses, thus helping address the problem of "Hallucination" Zhang et al. (2023) Tang and Yang (2024). However, naive RAG methods usually overlook the relationships among entities in the retrieved text chunks. To address this issue, RAG systems with graph structures were proposed Edge et al. (2024) Liang et al. (2024) Zhang et al. (2025) Peng et al. (2024b), which construct knowledge graphs (KGs) to model relationships between entities in the input documents. Although existing RAG systems integrating graph structures have demonstrated outstanding performance on various tasks, they still have some serious limitations. GraphRAG Edge et al. (2024) introduces communities in indexing using the Leiden algorithm Traag et al. (2019), but the communities only capture the structural proximity of the entities in the KG. KAG Liang et al. (2024) indexes with a hierarchical representation of information and knowledge, but their hierarchical structure relies too much on manual annotation and requires a lot of human domain knowledge, which renders their method not scalable to general tasks. LightRAG Guo et al. (2024) utilizes a dual-level retrieval approach to obtain local and global knowledge as the contexts for a query, but it ignores the knowledge gap between local and global knowledge, that is, local knowledge represented by the retrieved individual entities (i.e., entity-specific details) may not be semantically related to the global knowledge represented in the retrieved community summaries (i.e., broader, aggregated summaries), as these individual entities may not be a part of the retrieved communities for a query.
We highlight two critical challenges in existing RAG systems that integrate graph structures: (1) distant structural relationship between semantically similar entities and (2) knowledge gap between local and global knowledge. We illustrate them using a real example from a public dataset, as shown in Figure 1.
Challenge (1) occurs because existing methods over-rely on source documents, often resulting in constructing a knowledge graph (KG) with many entities that are not structurally proximate in the KG even though they share semantically similar attributes. For example, in Figure 1, although the entities "BIG DATA" and "RECOMMENDATION SYSTEM" share semantic relevance under the concept of "DATA MINING", their distant structural relationship in the KG reflects a corpus-driven disconnect. These inconsistencies between semantic relevance and structural proximity are systemic in KGs, undermining their utility in RAG systems where contextual coherence is critical.
Challenge (2) occurs as existing methods Guo et al. (2024) Edge et al. (2024) typically retrieve context either from global or local perspectives but fail to address the inherent disparity between these knowledge layers. Consider the query "Please introduce Amazon" in Figure 1, where global context emphasizes Amazon’s involvement in technological domains like big data and cloud computing, but local context retrieves entities directly linked to Amazon (e.g., subsidiaries, leadership). When these two knowledge layers are fed into LLMs as the contexts of a query without contextual alignment, LLMs may struggle to reconcile their distinct scopes, leading to disjointed reasoning, incomplete answers, or even contradictory outputs. For instance, an LLM might conflate Amazon’s role as a cloud provider (global) with its e-commerce operations (local), resulting in incoherent or factually inconsistent responses as the red words shown in the case. This underscores the need for new methods that bridge hierarchical knowledge layers to ensure cohesive reasoning in RAG systems.
To address these challenges, we propose Retrieval-Augmented Generation with Hierarchical Knowledge (HiRAG), which integrates hierarchical knowledge into the indexing and retrieval processes. Hierarchical knowledge Sarrafzadeh and Lank (2017) is a natural concept in both graph structure and human cognition, yet it has been overlooked in existing approaches. Specifically, to address Challenge (1), we introduce Indexing with Hierarchical Knowledge (HiIndex). Rather than simply constructing a flat KG, we index a KG hierarchically layer by layer. Each entity (or node) in a higher layer summarizes a cluster of entities in the lower layer, which can enhance the connectivity between semantically similar entities. For example, in Figure 1, the inclusion of the summary entity "DATA MINING" strengthens the connection between "BIG DATA" and "RECOMMENDATION SYSTEM". To address Challenge (2), we propose Retrieval with Hierarchical Knowledge (HiRetrieval). HiRetrieval effectively bridges local knowledge of entity descriptions to global knowledge of communities, thus resolving knowledge layer disparities. It provides a three-level context comprising the global level, the bridge level, and the local level knowledge to an LLM, enabling the LLM to generate more comprehensive and precise responses.
In summary, we make the following main contributions:
  • We identify and address two critical challenges in graph-based RAG systems: distant structural relationships between semantically similar entities and the knowledge gap between local and global information.
  • We propose HiRAG, which introduces unsupervised hierarchical indexing and a novel bridging mechanism for effective knowledge integration, significantly advancing the state-of-the-art in RAG systems.
  • Extensive experiments demonstrate both the effectiveness and efficiency of our approach, with comprehensive ablation studies validating the contribution of each component.
2 Related Work
In this section, we discuss recent research concerning graph-augmented LLMs, specifically RAG methods with graph structures. GNN-RAG Mavromatis and Karypis (2024) employs GNN-based reasoning to retrieve query-related entities. Then they find the shortest path between the retrieved entities and candidate answer entities to construct reasoning paths. LightRAG Guo et al. (2024) integrates a dual-level retrieval method with graph-enhanced text indexing. They also decrease the computational costs and speed up the adjustment process. GRAG Hu et al. (2024) leverages a soft pruning approach to minimize the influence of irrelevant entities in retrieved subgraphs. It also implements prompt tuning to help LLMs comprehend textual and topological information in subgraphs by incorporating graph soft prompts. StructRAG Li et al. (2024) identifies the most suitable structure for each task, transforms the initial documents into this organized structure, and subsequently generates responses according to the established structure. Microsoft GraphRAG Edge et al. (2024) first retrieves related communities and then let the LLM generate the response with the retrieved communities. They also answer a query with global search and local search. KAG Liang et al. (2024) introduces a professional domain knowledge service framework and employs knowledge alignment using conceptual semantic reasoning to mitigate the noise issue in OpenIE. KAG also constructs domain expert knowledge using human-annotated schemas. HippoRAG Gutiérrez et al. (2024) synergistically integrates LLMs, KGs, and Personalized PageRank to mimic the neocortex and hippocampus roles in human memory.
3 Preliminary and Definitions
In this section, we present the formulation of a graph-based RAG system, extending the definitions established in Guo et al. (2024) and  Peng et al. (2024a).
In a graph-based RAG system framework $\mathcal{M}$ as shown in Equation 1, $LLM$ is the generation module, $\mathcal{R}$ represents the retrieval module, $\varphi$ denotes the graph indexer, and $\psi$ refers to the graph retriever:
$\mathcal{M}=(LLM,\mathcal{R}(\varphi,\psi)).$ (1)
When we answer a query $q$, the graph-based RAG system produces an optimal response $a^{*}$ by maximizing a target distribution that captures the probability of generating each candidate response. Formally, this is defined as:
$a^{*}=arg\max_{a\in A}\mathcal{M}(a|q,\mathcal{G}),$ (2)
$\mathcal{G}=\varphi(\mathcal{D})=\{(h,r,t)|h,t\in\mathcal{V},r\in\mathcal{E}\},$ (3)
where $\mathcal{M}(a|q,\mathcal{G})$ is the target distribution with a graph retriever $\psi(G|q,\mathcal{G})$ and a generator $LLM(a|q,G)$, and $A$ is a set of possible responses. The graph database $\mathcal{G}$ is constructed from the original external database $\mathcal{D}$. We utilize the total probability formula to decompose $M(a|q,\mathcal{G})$, which can be expressed as
$\mathcal{M}(a|q,\mathcal{G})=\sum_{G\in\mathcal{G}}LLM(a|q,G)\cdot\psi(G|q,\mathcal{G}).$ (4)
Most of the time, we only need to retrieve the most relevant subgraph $G$ from the external graph database $\mathcal{G}$. Therefore, here we can approximate $\mathcal{M}(a|q,\mathcal{G})$ as follows:
$\mathcal{M}(a|q,\mathcal{G})\approx LLM(a|q,G^{*})\cdot\psi(G^{*}|q,\mathcal{G}),$ (5)
where $G^{*}$ denotes the optimal subgraph we retrieve from the external graph database $\mathcal{G}$. What we finally want is to get a better generated answer $a^{*}$.
Figure 2: The overall architecture of the HiRAG framework.
4 The HiRAG Framework
HiRAG consists of the two modules, HiIndex and HiRetrieval, as shown in Figure 2. In the HiIndex module, we construct a hierarchical KG with different knowledge granularity in different layers. The summary entities in a higher layer represent more coarse-grained, high-level knowledge but they can enhance the connectivity between semantically similar entities in a lower layer. In the HiRetrieval module, we select the most relevant entities from each retrieved community and find the shortest path to connect them, which serve as the bridge-level knowledge to connect the knowledge at both local and global levels. Then an LLM will generate responses with these three-level knowledge as the context.
4.1 Indexing with Hierarchical Knowledge
In the HiIndex module, we index the input documents as a hierarchical KG. First, we employ the entity-centric triple extraction to construct a basic KG $\mathcal{G}_{0}$ following Carta et al. (2023). Specifically, we split the input documents into text chunks with some overlaps. These chunks will be fed into the LLM with well-designed prompts to extract entities $\mathcal{V}_{0}$ first. Then the LLM will generate relations (or edges) $\mathcal{E}_{0}$ between pairs of the extracted entities based on the information of the corresponding text chunks. The basic KG can be represented as
$\mathcal{G}_{0}=\{(h,r,t)|h,t\in\mathcal{V}_{0},r\in\mathcal{E}_{0}\}.$ (6)
The basic KG is also the 0-th layer of our hierarchical KG. We denote the set of entities (nodes) in the $i$-th layer as $\mathcal{L}_{i}$ where $\mathcal{L}_{0}=\mathcal{V}_{0}$. To construct the $i$-th layer of the hierarchical KG, for $i\geq 1$, we first fetch the embeddings of the entities in the $(i-1)$-th layer of the hierarchical KG, which is denoted as
$\mathcal{Z}_{i-1}=\{Embedding(v)|v\in\mathcal{L}_{i-1}\},$ (7)
where $Embedding(v)$ is the embedding of an entity $v$. Then we employ Gaussian Mixture Models (GMMs) to conduct semantical clustering on $\mathcal{L}_{i-1}$ based on $\mathcal{Z}_{i-1}$, following the method described in RAPTOR Sarthi et al. (2024). We obtain a set of clusters as
$\mathcal{C}_{i-1}=GMM(\mathcal{L}_{i-1},\mathcal{Z}_{i-1})=\{\mathcal{S}_{1},\dots,\mathcal{S}_{c}\},$ (8)
where $\forall x,y\in[1,c]$, $|\mathcal{S}_{x}\cap\mathcal{S}_{y}|\geq 0$ and $\bigcup_{1\leq x\leq c}\mathcal{S}_{x}=\mathcal{L}_{i-1}$. After clustering with GMMs, the descriptions of the entities in each cluster in $\mathcal{C}_{i-1}$ are fed into the LLM to generate a set of summary entities for the $i$-th layer. Thus, the set of summary entities in the $i$-th layer, i.e., $\mathcal{L}_{i}$, is the union of the sets of summary entities generated from all clusters in $\mathcal{C}_{i-1}$. Then, we create the relations between entities in $\mathcal{L}_{i-1}$ and entities in $\mathcal{L}_{i}$, denoted as $\mathcal{E}_{\{i-1,i\}}$, by connecting the entities in each cluster $\mathcal{S}\in\mathcal{C}_{i-1}$ to the corresponding summary entities in $\mathcal{L}_{i}$ that are generated from the entities in $\mathcal{S}$.
To generate summary entities in $\mathcal{L}_{i}$, we use a set of meta summary entities $\mathcal{X}$ to guide the LLM to generate the summary entities. Here, $\mathcal{X}$ is a small set of general concepts such as "organization", "person", "location", "event", "technology", etc., that are generated by LLM. For example, the meta summary "technology" could guide the LLM to generate summary entities such as "big data" and "AI". Note that conceptually $\mathcal{X}$ is added as the top layer in Figure 2, but $\mathcal{X}$ is actually not part of the hierarchical KG.
After generating the summary entities and relations in the i-th layer, we update the KG as follows:
$\mathcal{E}_{i}=\mathcal{E}_{i-1}\cup\mathcal{E}_{\{i-1,i\}},$ (9)
$\mathcal{V}_{i}=\mathcal{V}_{i-1}\cup\mathcal{L}_{i},$ (10)
$\mathcal{G}_{i}=\{(h,r,t)|h,t\in\mathcal{V}_{i},r\in\mathcal{E}_{i}\}.$ (11)
We repeat the above process for each layer from the 1st layer to the $k$-th layer. We will discuss how to choose the parameter $k$ in Section 5. Also note that there is no relation between the summary entities in each layer except the 0-th layer (i.e., the basic KG).
We also employ the Leiden algorithm Traag et al. (2019) to compute a set of communities $\mathcal{P}$ from the hierarchical KG. Each community may contain entities from multiple layers and an entity may appear in multiple communities. For each community $p\in\mathcal{P}$, we generate an interpretable semantic report using LLMs. Unlike existing methods such as GraphRAG Edge et al. (2024) and LightRAG Guo et al. (2024), which identify communities based solely on direct structural proximity in a basic KG, our hierarchical KG introduces multi-resolution semantic aggregation. Higher-layer entities in our KG act as semantic hubs that abstract clusters of semantically related entities regardless of their distance from each other in a lower layer. For example, while a flat KG might separate "cardiologist" and "neurologist" nodes due to limited direct connections, their hierarchical abstraction as "medical specialists" in upper layers enables joint community membership. The hierarchical structure thus provides dual connectivity enhancement: structural cohesion through localized lower-layer connections and semantic bridging via higher-layer abstractions. This dual mechanism ensures our communities reflect both explicit relational patterns and implicit conceptual relationships, yielding more comprehensive knowledge groupings than structure-only approaches.
4.2 Retrieval with Hierarchical Knowledge
We now discuss how we retrieve hierarchical knowledge to address the knowledge gap issue. Based on the hierarchical KG $\mathcal{G}_{k}$ constructed in Section 4.1, we retrieve three-level knowledge at both local and global levels, as well as the bridging knowledge that connects them.
To retrieve local-level knowledge, we extract the top-$n$ most relevant entities $\hat{\mathcal{V}}$ as shown in Equation 12, where $Sim(q,v)$ is a function that measures the semantic similarity between a user query $q$ and an entity $v$ in the hierarchical KG $\mathcal{G}_{k}$. We set $n$ to 20 as default.
$\hat{\mathcal{V}}=TopN(\{v\in\mathcal{V}_{k}|Sim(q,v)\},n).$ (12)
To access global-level knowledge related to a query, we find the communities $\hat{\mathcal{P}}\subset\mathcal{P}$ that are connected to the retrieved entities as described in Equation 13, where $\mathcal{P}$ is computed during indexing in Section 4.1. Then the community reports of these communities are retrieved, which represent coarse-grained knowledge relevant to the user’s query.
$\hat{\mathcal{P}}=\bigcup_{p\in\mathcal{P}}\{p|p\cap\hat{\mathcal{V}}\neq\emptyset\}.$ (13)
To bridge the knowledge gap between the retrieved local-level and global-level knowledge, we also find a set of reasoning paths $\mathcal{R}$ connecting the retrieved communities. Specifically, from each community, we select the top-$m$ query-related key entities and collect them into $\hat{\mathcal{V}}_{\hat{\mathcal{P}}}$, as shown in Equation 14. The set of reasoning paths $\mathcal{R}$ is defined as the set of shortest paths between each pair of key entities according to their order in $\hat{\mathcal{V}}_{\hat{\mathcal{P}}}$, as shown in Equation 15. Based on $\mathcal{R}$, we construct a subgraph $\hat{\mathcal{R}}$ as described in Equation 16. Here, $\hat{\mathcal{R}}$ collects a set of triples from the KG that connect the knowledge in the local entities and the knowledge in the global communities.
$\hat{\mathcal{V}}_{\hat{\mathcal{P}}}=\bigcup_{p\in\hat{\mathcal{P}}}TopN(\{v\in p|Sim(q,v)\},m),$ (14)
$\mathcal{R}=\bigcup_{i\in[1,|\hat{\mathcal{V}}_{\hat{\mathcal{P}}}|-1]}ShortestPath_{\mathcal{G}_{k}}(\hat{\mathcal{V}}_{\hat{\mathcal{P}}}[i],\hat{\mathcal{V}}_{\hat{\mathcal{P}}}[i+1]),$ (15)
^=$\hat{\mathcal{R}}=\{(h,r,t)\in\mathcal{G}_{k}|h,t\in\mathcal{R}\}.$ (16)
After retrieving the three-level hierarchical knowledge, i.e., local-level descriptions of the individual entities in $\hat{\mathcal{V}}$, global-level community reports of the communities in $\hat{\mathcal{P}}$, and bridge-level descriptions of the triples in $\hat{\mathcal{R}}$, we feed them as the context to the LLM to generate a comprehensive answer to the query. We also provide the detailed procedures of HiRAG with pseudocodes in Appendix D.
4.3 Why is HiRAG effective?
HiRAG’s efficacy stems from its hierarchical architecture, HiIndex (i.e., hierarchical KG) and HiRetrieval (i.e., three-level knowledge retrieval), which directly mitigates the limitations outlined in Challenges (1) and (2) as described in Section 1.
Addressing Challenge (1): The hierarchical knowledge graph $\mathcal{G}_{k}$ introduces summary entities in its higher layers, creating shortcuts between entities that are distantly located in lower layers. This design bridges semantically related concepts efficiently, bypassing the need for exhaustive traversal of fine-grained relationships in the KG.
Resolving Challenge (2): HiRetrieval constructs reasoning paths by linking the top-$n$ entities most semantically relevant to a query with their associated communities. These paths represent the shortest connections between localized entity descriptions and global community-level insights, ensuring that both granular details and broader contextual knowledge inform the reasoning process.
Synthesis: By integrating (i) semantically similar entities via hierarchical shortcuts, (ii) global community contexts, and (iii) optimized pathways connecting local and global knowledge, HiRAG generates comprehensive, context-aware answers to user queries.
5 Experimental Evaluation
In this section. We report the performance evaluation results of HiRAG on Query-Focused Summarization (QFS) and Multi-Hop QA (MHQA) tasks respectively.
Baseline Methods. We compared HiRAG with state-of-the-art and popular baseline RAG methods, which are designed for QFS and MHQA tasks respectively. For QFS tasks: NaiveRAG Gao et al. (2022) Gao et al. (2023) splits original documents into chunks and retrieves relevant text chunks through vector search. GraphRAG Edge et al. (2024) utilizes communities and we use the local search mode in our experiments as it retrieves community reports as global knowledge, while their global search mode is known to be too costly and does not use local entity descriptions. LightRAG Guo et al. (2024) uses both global and local knowledge to answer a query. FastGraphRAG Circlemind (2024) integrates KG and personalized PageRank as proposed in HippoRAG Gutiérrez et al. (2024). KAG Liang et al. (2024) integrates structured reasoning of KG with LLMs and employs mutual indexing and logical-form-guided reasoning to enhance professional domain knowledge services. For MHQA tasks: in addition to NaiveRAG, GraphRAG, LightRAG, and FastGraphRAG, we also include RAPTOR, HippoRAG Gutiérrez et al. (2024) and HippoRAG2 Gutiérrez et al. (2025) as the state-of-the-art RAG methods for MHQA tasks.
Datasets and Queries. For QFS tasks, we used four datasets from the UltraDomain benchmark Qian et al. (2024), which is designed to evaluate RAG systems across diverse applications, focusing on long-context tasks and high-level queries in specialized domains. We used Mix, CS, Legal, and Agriculture datasets like in LightRAG Guo et al. (2024). We also used the benchmark queries provided in UltraDomain for each of the four datasets. For MHQA tasks, we randomly sampled 1,000 queries from 2WikiMultiHopQA Ho et al. (2020) and HotpotQA Yang et al. (2018) following the settings in HippoRAG2 Gutiérrez et al. (2025). The statistics of these datasets are given in Appendix B.
LLM. For QFS tasks, we employed DeepSeek-V3 DeepSeek-AI et al. (2024) as the LLM for information extraction, entity summarization, and answer generation in HiRAG and other baseline methods. We utilized GLM-4-Plus GLM et al. (2024) as the embedding model for vector search and semantic clustering because DeepSeek-V3 does not provide an accessible embedding model. For MHQA tasks, we used GPT-4o-mini and nvidia/NVEmbed-v2 Lee et al. (2024) as the LLM and embedding model used in HiRAG and other baseline methods respectively, following the same settings in HippoRAG2.
Table 1: Win rates (%) of HiRAG, its two variants (for ablation study), and baseline methods on QFS tasks.
Mix CS Legal Agriculture
NaiveRAG HiRAG NaiveRAG HiRAG NaiveRAG HiRAG NaiveRAG HiRAG
Comprehensiveness 16.6% 83.4% 30.0% 70.0% 32.5% 67.5% 34.0% 66.0%
Empowerment 11.6% 88.4% 29.0% 71.0% 25.0% 75.0% 31.0% 69.0%
Diversity 12.7% 87.3% 14.5% 85.5% 22.0% 78.0% 21.0% 79.0%
Overall 12.4% 87.6% 26.5% 73.5% 25.5% 74.5% 28.5% 71.5%
GraphRAG HiRAG GraphRAG HiRAG GraphRAG HiRAG GraphRAG HiRAG
Comprehensiveness 42.1% 57.9% 40.5% 59.5% 48.5% 51.5% 49.0% 51.0%
Empowerment 35.1% 64.9% 38.5% 61.5% 43.5% 56.5% 48.5% 51.5%
Diversity 40.5% 59.5% 30.5% 69.5% 47.0% 53.0% 45.5% 54.5%
Overall 35.9% 64.1% 36.0% 64.0% 45.5% 54.5% 46.0% 54.0%
LightRAG HiRAG LightRAG HiRAG LightRAG HiRAG LightRAG HiRAG
Comprehensiveness 36.8% 63.2% 44.5% 55.5% 49.0% 51.0% 38.5% 61.5%
Empowerment 34.9% 65.1% 41.5% 58.5% 43.5% 56.5% 36.5% 63.5%
Diversity 34.1% 65.9% 33.0% 67.0% 63.0% 37.0% 37.5% 62.5%
Overall 34.1% 65.9% 41.0% 59.0% 48.0% 52.0% 38.5% 61.5%
FastGraphRAG HiRAG FastGraphRAG HiRAG FastGraphRAG HiRAG FastGraphRAG HiRAG
Comprehensiveness 0.8% 99.2% 0.0% 100.0% 1.0% 99.0% 0.0% 100.0%
Empowerment 0.8% 99.2% 0.0% 100.0% 0.0% 100.0% 0.0% 100.0%
Diversity 0.8% 99.2% 0.5% 99.5% 1.5% 98.5% 0.0% 100.0%
Overall 0.8% 99.2% 0.0% 100.0% 0.0% 100.0% 0.0% 100.0%
KAG HiRAG KAG HiRAG KAG HiRAG KAG HiRAG
Comprehensiveness 2.3% 97.7% 1.0% 99.0% 16.5% 83.5% 5.0% 99.5%
Empowerment 3.5% 96.5% 4.5% 95.5% 9.0% 91.0% 5.0% 99.5%
Diversity 3.8% 96.2% 5.0% 95.0% 11.0% 89.0% 3.5% 96.5%
Overall 2.3% 97.7% 1.5% 98.5% 8.5% 91.5% 0.0% 100.0%
w/o HiIndex HiRAG w/o HiIndex HiRAG w/o HiIndex HiRAG w/o HiIndex HiRAG
Comprehensiveness 46.7% 53.3% 44.2% 55.8% 49.0% 51.0% 50.5% 49.5%
Empowerment 43.2% 56.8% 38.8% 61.2% 47.5% 52.5% 50.5% 49.5%
Diversity 40.5% 59.5% 40.0% 60.0% 48.0% 52.0% 48.5% 51.5%
Overall 42.4% 57.6% 40.0% 60.0% 46.5% 53.5% 48.0% 52.0%
w/o Bridge HiRAG w/o Bridge HiRAG w/o Bridge HiRAG w/o Bridge HiRAG
Comprehensiveness 49.2% 50.8% 46.5% 53.5% 49.5% 50.5% 47.0% 53.0%
Empowerment 44.2% 55.8% 43.0% 57.0% 38.5% 61.5% 41.0% 59.0%
Diversity 44.6% 55.4% 44.0% 56.0% 43.5% 56.5% 46.0% 54.0%
Overall 47.3% 52.7% 42.5% 57.5% 44.0% 56.0% 42.0% 58.0%
Table 2: EM and F1 scores (%) of HiRAG, and baseline methods on MHQA tasks. The best and second-best results are highlighted
2WikiMultiHopQA HotpotQA Average
Method EM F1 EM F1 EM F1
NaiveRAG 54.4 60.8 57.3 71.0 55.9 65.9
RAPTOR 39.7 48.4 50.6 64.7 45.2 56.6
GraphRAG 45.7 61.0 51.4 67.6 48.6 64.3
LightRAG 2.5 12.1 9.9 20.2 6.3 16.2
FastGraphRAG 20.8 44.8 35.0 49.6 27.9 47.2
HippoRAG 59.4 67.3 46.3 60.0 52.9 63.7
HippoRAG2 60.5 69.7 56.3 71.1 58.4 70.4
HiRAG 69.0 74.4 62.0 72.9 65.5 73.7
5.1 Overall Performance Comparison
QFS Task Evaluation Details. For QFS tasks, our experiments followed the evaluation methods of recent work Edge et al. (2024)Guo et al. (2024) by employing a powerful LLM to conduct multi-dimensional comparison. We used the win rate to compare different methods, which indicates the percentage of instances that a method generates higher-quality answers compared to another method as judged by the LLM. We utilized GPT-4o Achiam et al. (2023) as the evaluation model to judge which method generates a superior answer for each query for the following four dimensions: (1) Comprehensiveness: how thoroughly does the answer address the question, covering all relevant aspects and details? (2) Empowerment: how effectively does the answer provide actionable insights or solutions that empower the user to take meaningful steps? (3) Diversity: how well does the answer incorporate a variety of perspectives, approaches, or solutions to the problem? (4) Overall: how does the answer perform overall, considering comprehensiveness, empowerment, diversity, and any other relevant factors? For a fair comparison, we also alternated the order of the answers generated by each pair of methods in the prompts and calculated the overall win rates of each method. We also report the reasonability and credibility of using a powerful LLM as a judge in Appendix G.
MHQA Task Evaluation Details. For MHQA tasks, our experiments also followed the evaluation methods of recent work Gutiérrez et al. (2024) Gutiérrez et al. (2025) by leveraging two established metrics: Exact Match (EM) and F1 scores, which are applied to the generated answers. Compared with the metric of win rates used in QFS tasks, the performance with EM and F1 scores can indicate HiRAG’s ability to achieve correctness.
Evaluation Results. We present the win rates, EM, and F1 scores of HiRAG and various baseline methods in Table 1 and Table 2. HiRAG consistently outperforms existing approaches in both QFS tasks across all four datasets and four evaluation dimensions and MHQA tasks across all two datasets in the majority of cases. Key insights derived from the results are summarized below.
Graph structure enhances RAG systems: NaiveRAG exhibits inferior performance compared to methods integrating graph structures, primarily due to its inability to model relationships between entities in retrieved components. Furthermore, its context processing is constrained by the token limitations of LLMs, highlighting the importance of structured knowledge representation for robust retrieval and reasoning.
Global knowledge improves answer quality: Approaches incorporating global knowledge (GraphRAG, LightRAG, KAG, HiRAG) significantly surpass FastGraphRAG, which relies on local knowledge via personalized PageRank. Answers generated without global context lack depth and diversity, underscoring the necessity of holistic knowledge integration for comprehensive responses.
HiRAG’s superior performance: Among graph-enhanced RAG systems, HiRAG achieves the highest performance across all datasets (spanning diverse domains) and evaluation dimensions. This superiority stems primarily from two innovations: (1) HiIndex which enhances connections between remote but semantically similar entities in the hierarchical KG, and (2) HiRetrieval which effectively bridges global knowledge with localized context to optimize relevance and coherence.
5.2 Hierarchical KG vs. Flat KG
To evaluate the significance of the hierarchical KG, we replace the hierarchical KG with a flat KG (or a basic KG), denoted by w/o HiIndex as reported in Table 1. Compared with HiRAG, the win rates of w/o HiIndex drop in almost all cases and quite significantly in at least half of the cases. This ablation study thus shows that the hierarchical indexing plays an important role in the quality of answer generation, since the connectivity among semantically similar entities is enhanced with the hierarchical KG, with which related entities can be grouped together both from structural and semantical perspectives.
From Table 1, we also observe that the win rates of w/o HiIndex are better or comparable to those of GraphRAG and LightRAG when compared with HiRAG. This suggests that our three-level knowledge retrieval method, i.e., HiRetrieval, is effective even applied on a flat KG, because GraphRAG and LightRAG also index on a flat KG but they only use the local entity descriptions and global community reports, while w/o HiIndex uses an additional bridge-level knowledge.
Figure 3: Answer to the query in Figure 1 with additional bridge-level knowledge.
5.3 HiRetrieval vs. Gapped Knowledge
To show the effectiveness of HiRetrieval, we also created another variant of HiRAG without using the bridge-level knowledge, denoted by w/o Bridge in Table 1. The result shows that without the bridge-layer knowledge, the win rates drop significantly across all datasets and evaluation dimensions, because there is knowledge gap between the local-level and global-level knowledge as discussed in Section 1. We also report the knowledge coverage of bridge-level descriptions in Appendix H, which further proofs both local- and global-level knowledge are well connected in the bridge-level descriptions.
Case Study. Figure 3 shows the three-level knowledge used as the context to an LLM to answer the query in Figure 1. The bridge-level knowledge contains entity descriptions from different communities, as shown by the different colors in Figure 3, which helps the LLM correctly answer the question about Amazon’s role as an e-commerce and cloud provider.
5.4 Determining the Number of Layers
Figure 4: Cluster sparsity $CS_{i}$ and change rate from $CS_{i}$ to $CS_{i+1}$, where the shadow areas represent the value ranges of the four datasets and the blue/pink lines are the respective average values.
One important thing in HiIndex is to determine the number of layers, $k$, for the hierarchical KG, which should be determined dynamically according to the quality of clusters in each layer. We stop building another layer when the majority of the clusters consist of only a small number of entities, meaning that the entities can no longer be effectively grouped together. To measure that, we introduce the notion of cluster sparsity $CS_{i}$, as inspired by graph sparsity, to measure the quality of clusters in the $i$-th layer as described in Equation 17.
$CS_{i}=1-\frac{\sum_{\mathcal{S}\in\mathcal{C}_{i}}|\mathcal{S}|(|\mathcal{S}|-1)}{|\mathcal{L}_{i}|(|\mathcal{L}_{i}|-1)}.$ (17)
The more the clusters in $\mathcal{C}_{i}$ have a small number of entities, the larger is $CS_{i}$, where the worst case is when each cluster contains only one entity (i.e., $CS_{i}=1$). Figure 4 shows that as we have more layers, the cluster sparsity increases and then stabilizes. We also plot the change rate from $CS_{i}$ to $CS_{i+1}$, which shows that there is little or no more change after constructing a certain number of layers. We set a threshold $\epsilon=5\%$ and stop constructing another layer when the change rate of cluster sparsity is lower than $\epsilon$ because the cluster quality has little or no improvement after that.
5.5 Efficiency and Costs Analysis
To evaluate the efficiency and costs of HiRAG, we also report the token costs, the number of API calls, and the time costs of indexing and retrieval of HiRAG and the baselines in Table 5 in Appendix C. Although HiRAG needs more time and resources to conduct indexing for better performance, we remark that indexing is offline and the total cost is only about 7.55 USD for the Mix dataset using DeepSeek-V3. In terms of retrieval, unlike KAG and LightRAG, HiRAG does not cost any tokens for retrieval. Therefore, HiRAG is more efficient for online retrieval.
6 Conclusions
We presented a new approach to enhance RAG systems by effectively utilizing graph structures with hierarchical knowledge. By developing (1) HiIndex which enhances structural and semantic connectivity across hierarchical layers, and (2) HiRetrieval which effectively bridges global conceptual abstractions with localized entity descriptions, HiRAG achieves superior performance than existing methods.
Appendix
Appendix A Limitations
HiRAG has the following limitations. Firstly, constructing a high-quality hierarchical KG may incur substantial token consumption and time overhead, as LLMs need to perform entity summarization in each layer. However, the monetary cost of using LLMs may not be the major concern as the cost is decreasing rapidly recently, and therefore we may consider parallelizing the indexing process to reduce the indexing time. Secondly, the retrieval module requires more sophisticated query-aware ranking mechanisms. Currently, our HiRetrieval module relies solely on LLM-generated weights for relation ranking, which may affect query relevance. We will research for more effective ranking mechanisms to further improve the retrieval quality.
Appendix B Experimental Datasets
Table 3: Statistics of QFS task datasets.
Dataset Mix CS Legal Agriculture
# of Documents 61 10 94 12
# of Tokens 625,948 2,210,894 5279400 2,028,496
Table 4: Statistics of MHQA task datasets.
Dataset 2WikiMultiHopQA HotpotQA
# of Queries 1,000 1,000
# of Passages 6,119 9,811
# of Tokens 1,189,866 1,507,487
Table 3 and Table 4 presents the statistical characteristics of the experimental datasets, where all documents were consistently tokenized using Byte Pair Encoding (BPE) tokenizer "cl100k_base".
Appendix C Cost Analysis
Table 5: Comparisons in terms of tokens, API calls and time cost across four datasets.
Token Cost API Calls Time Cost (s)
Dataset Method Indexing Retrieval Indexing Retrieval Indexing Retrieval
Mix GraphRAG 8,507,697 0.00 2,666 1.00 6,696 0.70
LightRAG 3,849,030 357.76 1,160 2.00 3,342 3.06
KAG 6,440,668 110,532.00 831 9.17 8,530 58.47
HiRAG 21,898,765 0.00 6,790 1.00 17,208 0.85
CS GraphRAG 27,506,689 0.00 8,649 1.00 19,255 0.98
LightRAG 12,638,997 353.37 3,799 2.00 14,307 4.97
KAG 7,358,717 89,746.00 2,190 6.29 14,837 46.37
HiRAG 56,042,906 0.00 16,535 1.00 44,994 1.17
Legal GraphRAG 51,168,359 0.00 13,560 1.00 30,065 1.12
LightRAG 30,299,958 353.77 9,442 2.00 21,505 5.44
KAG 18,431,706 97,683.00 4,980 7.82 29,191 51.26
HiRAG 106,427,778 0.00 27,224 1.00 115,232 2.04
Agriculture GraphRAG 27,974,472 0.00 8,669 1.00 20,362 1.17
LightRAG 12,031,096 354.62 3,694 2.00 13,550 5.64
KAG 7,513,424 93,217.00 2,358 6.83 22,557 49.57
HiRAG 96,080,883 0.00 22,736 1.00 50,920 1.76
As shown in Table 5, we calculate the average token, API calls and time costs of different methods across four datasets. For indexing, we record the total costs of the entire indexing process. For retrieval, we calculate the average costs per query during the retrieval process, which could reflect the performance while the methods are deployed online.
0:  Basic knowledge graph $\mathcal{G}_{0}$ extracted by the LLM; Predefined threshold $\epsilon$;
0:  Hierarchical knowledge graph $\mathcal{G}_{k}$;
1:  $\mathcal{L}_{0}\leftarrow\mathcal{V}_{0}$;
2:  $\mathcal{Z}_{0}\leftarrow\{Embedding(v)|v\in\mathcal{L}_{0}\}$;
3:  $i\leftarrow 1$;
4:  while True do
5:     /*Perform semantical clustering*/
6:     $\mathcal{C}_{i-1}\leftarrow GMM(\mathcal{G}_{i-1},\mathcal{Z}_{i-1})$;
7:     /*Calculate cluster sparsity*/
8:     $CS_{i}\leftarrow 1-\frac{\sum_{\mathcal{S}\in\mathcal{C}_{i-1}}|\mathcal{S}|(|\mathcal{S}|-1)}{|\mathcal{L}_{i-1}|(|\mathcal{L}_{i-1}|-1)}$;
9:     if change rate of $CS_{i}$ $\leq\epsilon$ then
10:        $i\leftarrow i-1$;
11:        break;
12:     end if
13:     /*Generate summary entities and relations*/
14:     $\mathcal{L}_{i}\leftarrow\{\}$;
15:     $\mathcal{E}_{\{i-1,i\}}\leftarrow\{\}$;
16:     for $\mathcal{S}_{x}$ in $\mathcal{C}_{i-1}$ do
17:        $\mathcal{L},\mathcal{E}\leftarrow LLM(\mathcal{S}_{x},\mathcal{X})$;
18:        $\mathcal{L}_{i}\leftarrow\mathcal{L}_{i}\cup\mathcal{L}$;
19:        $\mathcal{E}_{\{i-1,i\}}\leftarrow\mathcal{E}_{\{i-1,i\}}\cup\mathcal{E}$;
20:     end for
21:     $\mathcal{Z}_{i}=\{Embedding(v)|v\in\mathcal{L}_{i}\}$;
22:     /*Update KG*/
23:     $\mathcal{E}_{i}\leftarrow\mathcal{E}_{i-1}\cup\mathcal{E}_{\{i-1,i\}}$;
24:     $\mathcal{V}_{i}\leftarrow\mathcal{V}_{i-1}\cup\mathcal{L}_{i}$;
25:     $\mathcal{G}_{i}\leftarrow\{(h,r,t)|h,t\in\mathcal{V}_{i},r\in\mathcal{E}_{i}\}$
26:     $i\leftarrow i+1$;
27:  end while
28:  $k\leftarrow i$;
29:  $\mathcal{G}_{k}\leftarrow\{(h,r,t)|h,t\in\mathcal{V}_{k},r\in\mathcal{E}_{k}\}$;
Algorithm 1 HiIndex
0:  The hierarchical knowledge graph $\mathcal{G}_{k}$; The detected community set $\mathcal{P}$ in $\mathcal{G}_{k}$; The number of retrieved entities $n$; The number of selected key entities $m$ in each retrieved community;
0:  The generated answer $a$;
1:  /*The local-layer knowledge context*/
2:  $\hat{\mathcal{V}}\leftarrow TopN(\{v\in\mathcal{V}_{k}|Sim(v,q)\},n)$;
3:  /*The global-layer knowledge context*/
4:  $\hat{\mathcal{P}}\leftarrow\bigcup_{p\in\mathcal{P}}\{p|p\cap\hat{\mathcal{V}}\neq\phi\}$;
5:  $^{}^\hat{\mathcal{R}}\leftarrow\{\}$;
6:  $\hat{\mathcal{V}}_{\hat{\mathcal{P}}}\leftarrow\{\}$;
7:  /*Select key entities*/
8:  for $p$ in $\hat{\mathcal{P}}$ do
9:     $\hat{\mathcal{V}}_{\hat{\mathcal{P}}}\leftarrow\hat{\mathcal{V}}_{\hat{\mathcal{P}}}\cup TopN(\{v\in p|Sim(v,q)\},m)$;
10:  end for
11:  /*Find the reasoning path*/
12:  for $i$ in $[1,|\hat{\mathcal{V}}_{\hat{\mathcal{P}}}|-1]$ do
13:     $\mathcal{R}\leftarrow\mathcal{R}\cup ShortestPath_{\mathcal{G}_{k}}(\hat{\mathcal{V}}_{\hat{\mathcal{P}}}[i],\hat{\mathcal{V}}_{\hat{\mathcal{P}}}[i+1])$;
14:  end for
15:  /*The bridge-layer knowledge context*/
16:  ^$\hat{\mathcal{R}}\leftarrow\{(h,r,t)\in\mathcal{G}_{k}|h,t\in\mathcal{R}\}$;
17:  /*Generate the answer*/
18:  $a\leftarrow LLM(q,\hat{\mathcal{V}},\hat{R},\hat{\mathcal{P}})$;
Algorithm 2 HiRetrieval
Appendix D Implementation Details of HiRAG
We give a more detailed and formulated expression of hierarchical indexing (HiIndex) and hierarchical retrieval (HiRetrieval). As described in Algorithm 1, the hierarchical knowledge graph is constructed iteratively. The number of clustered layers depends on the rate of change in the cluster sparsity at each layer. As shown in Algorithm 2, we retrieve knowledge of three layers (local layer, global layer, and bridge layer) as contexts for LLM to generate more comprehensive and accurate answers.
Figure 5: Comparisons between the clustering coefficients of GraphRAG, LightRAG and HiRAG across four datasets.
Appendix E The Clustering Coefficients of HiIndex
We calculate and compare the clustering coefficients of GraphRAG, LightRAG and HiRAG in Figure 5. It evaluates the average connectivity among the neighbors of each entity of our hierarchical KG using the global clustering coefficient
$C=\frac{closed\_triples}{all\_triples},$ (18)
which represents how frequently entities form triangles (closed triples). As shown in Figure 5, HiRAG demonstrates a higher clustering coefficient than other baseline methods, which means that more entities in the hierarchical KG constructed by the HiIndex module tend to cluster together. And the hierarchical KG maintains a proper hierarchy without becoming an overly connected clique (i.e., C=1) or a near-clique. These are the reasons why the HiIndex module can improve the performance of RAG systems.
Figure 6: The shortest path with hierarchical KG between the entities in the case mentioned in the introduction.
Appendix F A Simple Case of Hierarchical KG
As shown in Figure 6, we fix the issues mentioned in Section 1 with a hierarchical KG. This case demonstrates that the GMMs clustered semantically similar entities "BIG DATA" and "RECOMMENDATION SYSTEM" together. The LLM summarizes "DISTRIBUTED COMPUTING" as their shared summary entities in the next layer. As a consequence, the connections between these related entities can be enhanced from a semantic perspective.
Table 6: Win rates (%) of HiRAG and baseline methods on QFS tasks with three different powerful LLMs as the evaluator.
GPT-4o Claude-3.5-sonnet Qwen-turbo
NaiveRAG HiRAG NaiveRAG HiRAG NaiveRAG HiRAG
Comprehensiveness 16.6% 83.4% 13.0% 87.0% 13.6% 86.4%
Empowerment 11.6% 88.4% 10.0% 90.0% 12.7% 87.3%
Diversity 12.7% 87.3% 28.0% 72.0% 18.2% 81.8%
Overall 12.4% 87.6% 11.0% 89.0% 12.7% 87.3%
GraphRAG HiRAG GraphRAG HiRAG GraphRAG HiRAG
Comprehensiveness 42.1% 57.9% 39.1% 60.9% 32.4% 67.6%
Empowerment 35.1% 64.9% 31.8% 68.2% 33.3% 66.7%
Diversity 40.5% 59.5% 48.2% 51.8% 40.7% 59.3%
Overall 35.9% 64.1% 32.7% 67.3% 32.4% 67.6%
LightRAG HiRAG LightRAG HiRAG LightRAG HiRAG
Comprehensiveness 36.8% 63.2% 36.4% 63.6% 35.5% 64.5%
Empowerment 34.9% 65.1% 31.8% 68.2% 35.5% 64.5%
Diversity 34.1% 65.9% 40.1% 59.1% 39.1% 60.9%
Overall 34.1% 65.9% 33.6% 66.4% 35.5% 64.5%
FastGraphRAG HiRAG FastGraphRAG HiRAG FastGraphRAG HiRAG
Comprehensiveness 0.8% 99.2% 0.0% 100.0% 0.8% 99.2%
Empowerment 0.8% 99.2% 0.0% 100.0% 0.8% 99.2%
Diversity 0.8% 99.2% 0.9% 99.1% 0.8% 99.2%
Overall 0.8% 99.2% 0.0% 100.0% 0.8% 99.2%
KAG HiRAG KAG HiRAG KAG HiRAG
Comprehensiveness 2.3% 97.7% 1.8% 98.2% 3.6% 96.4%
Empowerment 3.5% 96.5% 2.7% 97.3% 5.5% 94.5%
Diversity 3.8% 96.2% 12.7% 87.3% 10.9% 89.1%
Overall 2.3% 97.7% 1.8% 98.2% 3.6% 96.4%
Appendix G Cross Validation for LLM as a Judge
Although our LLM-based evaluation approach for QFS tasks is a common practice in the performance evaluation by existing graph RAG methods Es et al. [2024] Guo et al. [2024] Edge et al. [2024], we also conducted cross-verification using Qwen-turbo and Claude-3.5-sonnet as the LLM judge to further make our experimental results more convincing. As shown in Table 6, the results consistently demonstrate HiRAG’s superiority over all the other graph RAG methods compared, confirming that our conclusions remain stable even when neutralizing LLM evaluator-specific biases. To reduce the cost, we report the results on the Mix dataset while the results on the other datasets follow a similar pattern.
Table 7: The average knowledge coverage of bridge-level descriptions across four QFS datasets.
Mix CS Legal Agriculture
Recall (%) Recall (%) Recall (%) Recall (%)
Global-level Info 38.61 48.96 53.44 50.75
Local-level Info 83.07 81.88 78.13 64.47
Appendix H Knowledge Coverage of Bridge-Level Descriptions
To further validate that both local- and global-level knowledge are well connected in the bridge-level descriptions., we counted the average token-level recall ratio of local- and global-level context in the bridge-level context across four QFS datasets. As shown in Table 7, the results demonstrate that our bridge-level retrieval effectively captures a significant portion of both entity-level and community-level information, providing empirical support for the method’s effectiveness. Here, the recall indicates that the percentage of global-level info or local-level info that is captured in the bridge-level context.
Appendix I Prompt Templates used in HiRAG
Figure 7: The prompt template designed to extract entities from text chunks.
I.1 Prompt Templates for Entity Extraction
As shown in Figure 7, we used that prompt template to extract entities from text chunks. We also give three examples to guide the LLM to extract entities with higher accuracy.
I.2 Prompt Templates for Relation Extraction
As shown in Figure 8, we extract relations from the entities extracted earlier and the corresponding text chunks. Then we can get the triples in the basic knowledge graph, which is also the 0-th layer of the hierarchical knowledge graph.
I.3 Prompt Templates for Entity Summarization
As shown in Figure 9, we generate summary entities in each layer of the hierarchical knowledge graph. We will not only let the LLM generate the summary entities from the previous layer, but also let it generate the relations between the entities of these two layers. These relations will clarify the reasons for summarizing these entities.
I.4 Prompt Templates for RAG Evaluation
In terms of the prompt templates we use to conduct evaluations, we utilize the same prompt design as that in LightRAG. The prompt will let the LLM generate both evaluation results and the reasons in JSON format to ensure clarity and accuracy.
Figure 8: The prompt template designed to extract relations from entities and text chunks.
Figure 9: The prompt template designed to generate summary entities and the corresponding relations.